given cos(A-B)=2cos(A+B) show that 3 tan(a) =1/tan(b)

Grade:10

2 Answers

Hend
6 years ago
Cos(A-B)=2cos(A+B)cosAcosB+sinAsinB=2(cosAcosB-sinAsinB)cosAcosB+sinAsinB————————— = 2cosAcosB-sinAsinB(÷everything by cosAcosB)1+tanAtanB—————— = 21–tanAtanB1+tanAtanB=2–2tanAtanBtanAtanB+2tanAtanB=2–13tanAtanB=1››3tanA= 1 ––––– tanB
Hend
6 years ago
Sorry when I typed it out the lines were supposed to mean division and I clicked enter but when I submitted my answer, it got messed up so just ignore it. Simply manipulate what is given to you in the log tables and its basic algebra from there.